# Machine Translation Models are Zero-Shot Detectors of Translation Direction

### v1

@@ 1 Introduction
While the original translation direction of parallel text is often ignored or unknown in the machine translation community, research has shown that it can be relevant for training Kurokawa et al. (2009); Ni et al. (2022) and evaluation Graham et al. (2020).222As of today, training data is not typically filtered by translation direction, but we find evidence of a need for better detection in recent work. For example, Post and Junczys-Dowmunt (2023) show that back-translated data is more suited than crawled parallel data for document-level training, presumably because of translations in the crawled data that lack document-level consistency.
Beyond machine translation, translation direction detection has practical applications in areas such as forensic linguistics, where determining the original of a document pair may help resolve plagiarism or forgery accusations. This study is partially inspired by a highly publicized criminal case in 2022, where one party has been accused of plagiarizing their (German) PhD thesis from an ostensibly older English book, but where the English book is suspected to be a forgery and translation of the thesis, possibly created with the explicit purpose of slandering the author of the thesis Ebbinghaus (2022); Zenthöfer (2022); Wikipedia (2023).
@@
In this work, we explore the unsupervised detection of translation directions purely on the basis of a neural machine translation (NMT) system’s translation probabilities in both directions, hypothesizing that . If the translation has been automatically generated, this hypothesis can be motivated by the fact that machine translation systems typically generate text with mode-seeking search algorithms, and consequently tend to over-produce high-frequency outputs and reduce lexical diversity Vanmassenhove et al. (2019). However, even human translations are known for so-called translationese properties such as interference, normalization, and simplification, and a (relative) lack of lexical diversity Teich (2003); Volansky et al. (2013); Toral (2019). We hypothesize that our detection method can be applied to both human and automatic translations.
We experiment with three different multilingual machine translation (MMT) systems to generate translation probabilities and test the approach on 20 translation directions with different resource availability. Our experiments show that our approach is effective, detecting the correct translation direction of human translations with an accuracy of 66% on average on the sentence-level, and 80% for documents with 10 sentences. For the output of neural MT systems, detection accuracy is even higher, but our hypothesis that does not hold for the output of pre-neural systems.
In a qualitative analysis, we show that the detection performance can be connected to the translationese characteristics found in the sentences. Finally, we apply the detection method to data from the aforementioned plagiarism allegation case. Our results provide further evidence suggesting that the document written in German is likely the original.
@@ 3 Methods
Given a parallel sentence pair , the main task in this work is to identify the translation direction between a language X and a language Y, and, consequently, establish which side is the original and which the translation. This is achieved by comparing the conditional translation probability by an NMT model with the conditional translation probability by a model operating in the inverse direction. Our core assumption is that translations are assigned higher conditional probabilities than the original by NMT models, so if , we predict that is the translation, and the original, and the original translation direction is . Figure 1 illustrates this process.
@@ 3.1 Detection on the Sentence Level
We follow earlier work by Junczys-Dowmunt (2018); Thompson and Post (2020), and average token-level (log-)probabilities.333Note that .
@@ 4 Experiments: Models and Data
We experiment with three massively multilingual machine translation models: M2M-100-418M (Fan et al., 2021), SMaLL-100 (Mohammadshahi et al., 2022), and NLLB-200-1.3B (Team et al., 2022).
@@
| Total | |||||
As an additional “real-world” dataset, we use 86 parallel sentences in German and English from a publicly documented plagiarism allegation case, in which translation-based plagiarism was the main focus (Ebbinghaus, 2022; Zenthöfer, 2022). The sentences were extracted from aligned excerpts of both the PhD thesis and the alleged source that were presented in a plagiarism analysis report (Weber, 2022). We extracted the sentences with OCR and manually checked for OCR mistakes. This dataset introduces an element of real-world complexity, as it involves translations that might not adhere strictly to professional standards nor does it have to fit into either of the aforementioned categories of translation strategy. It also provides a unique opportunity to test the robustness and adaptability of the translation direction detection system in a scenario that extends beyond controlled environments.
@@ 5.1 Sentence-level Classification
| Language Pair | Percentage of | Percentage of |
| HT bnhi | 66.20 | 32.61 |
| HT csuk | 42.69 | 57.31 |
| HT defr | 84.78 | 15.22 |
| HT xhzu | 46.12 | 55.99 |
A second result is that the translation detection works best for neural translations (75.0% macro-average), second-best for human translations (66.5% macro-average), and worst for pre-neural (41.5% macro-average). The fact that performance for pre-neural systems is below chance level indicates that our MMT models tend to assign low probabilities to the (often ungrammatical) outputs of pre-neural systems. In practice, one could pair our detection method with a monolingual model to identify such low-quality outputs.
@@ 5.4 Qualitative Analysis
Sentences |
Rel. Difference | |||
| 1 | DE: Mit dem Programm "Guten Tag, liebes Glück" ist er seit 2020 auf Tour. | |||
EN: He has been on tour with the programme "Guten Tag, liebes Glück" since 2020. (HT) |
0.145 | 0.558 | 0.26 | |
EN: He has been on tour since 2020. (NMT) |
0.272 | 0.092 | 2.95 | |
| 2 | EN: please try to perfprm thsi procedures" | |||
DE: bitte versuchen Sie es mit diesen Verfahren (HT) |
0.246 | 0.010 | 24.29 | |
DE: Bitte versuchen Sie, diese Prozeduren durchzuführen" (NMT) |
0.586 | 0.025 | 23.59 | |
| 3 | EN: If costs for your country are not listed, please contact us for a quote. | |||
DE: Wenn die Kosten für Ihr Land nicht aufgeführt sind, wenden Sie sich für einen Kostenvoranschlag an uns. (HT) |
0.405 | 0.525 | 0.77 | |
DE: Wenn die Kosten für Ihr Land nicht aufgeführt sind, kontaktieren Sie uns bitte für ein Angebot. (NMT) |
0.697 | 0.585 | 1.19 | |
| 4 | EN: Needless to say, it was chaos. | |||
DE: Es war natürlich ein Chaos. (HT) |
0.119 | 0.372 | 0.32 | |
DE: Unnötig zu sagen, es war Chaos. (NMT) |
0.755 | 0.591 | 1.28 | |
| 5 | DE: Mit freundlichen Grüßen | |||
FR: Cordialement (HT) |
0.026 | 0.107 | 0.24 | |
FR: Sincèrement (NMT) |
0.015 | 0.083 | 0.18 | |
FR: Sincères amitiés (NMT) |
0.062 | 0.160 | 0.39 | |
FR: Avec mes meilleures salutations (NMT) |
0.215 | 0.353 | 0.61 |
@@
Misclassified short sentences as in Example 5 are not a rarity in our experiments. Our findings show that reliable detection of translation direction, with an average accuracy exceeding 50%, is consistently attained for all language pairs we tested starting at a sentence length between 50 and 60 characters.555We used SMaLL-100 for this analysis. Additionally, we observed a trend where the accuracy of direction detection increases as the length of the sentences grows. This aligns with previous unsupervised approaches, which also documented higher accuracy the larger the text chunks, that were used, although there, reliable results were reported on a more extreme scale starting from text chunks with a length of 250 tokens Rabinovich and Wintner (2015).
@@
| DEEN | ENDE | |
EN: After 30 sec, the trypsin-containing PBS was poured off, and the cells were placed in the incubator for another minute. |
0.442 | 0.285 |
5.5 Document-Level Classification
Accuracy scores for document-level results by M2M-100 (best-performing system at sentence level) for high-resource language pairs are presented in Tables 8 (NMT) and 7 (HT). We consider documents with at least 10 sentences, and language pairs with at least 100 such documents in both directions.
5.6 Hypothesis Test on Real-World Dataset
@@
Overall, our analysis supports the hypothesis that German is indeed the language of origin in this real-world dataset. Nevertheless, we recommend that additional evidence be considered before drawing a final conclusion, given the error rate of 5–21% that we observed in our earlier experiments on English–German WMT documents.
@@ 6 Conclusion
We proposed a novel approach to detecting the translation direction of parallel sentences, using only an off-the-shelf multilingual NMT system. Experiments on WMT data showed that our approach, without any task-specific supervision, is able to detect the translation direction of NMT-translated sentences with relatively high accuracy. The accuracy increases to 96% if the classifier is provided with at least 10 sentences per document. We also found a robust accuracy for translations by human translators. Finally, we applied our approach to a real-world plagiarism allegation case and found that it supports the hypothesis that German is the original language in this case.
@@ Appendix B Comparison of Models (Document Level)
Appendix C Data Statistics

### v2

@@ 1 Introduction
While the original translation direction of parallel text is often ignored or unknown in the machine translation community, research has shown that it can be relevant for training Kurokawa et al. (2009); Ni et al. (2022) and evaluation Graham et al. (2020).222As of today, training data is not typically filtered by translation direction, but we find evidence of a need for better detection in recent work. For example, Post and Junczys-Dowmunt (2023) show that back-translated data is more suited than crawled parallel data for document-level training, presumably because of translations in the crawled data that lack document-level consistency. Beyond machine translation, translation direction detection has practical applications in areas such as forensic linguistics, where determining the original of a document pair may help resolve plagiarism or forgery accusations.
@@
In this work, we explore the unsupervised detection of translation directions purely on the basis of a neural machine translation (NMT) system’s translation probabilities in both directions. As illustrated in Figure 1, we hypothesize that , which, if it generally holds, would allow us to infer the original translation direction.
If the translation has been automatically generated, this hypothesis can be motivated by the fact that machine translation systems typically generate text with mode-seeking search algorithms, and consequently tend to over-produce high-frequency outputs and reduce lexical diversity Vanmassenhove et al. (2019). However, even human translations are known for so-called translationese properties such as interference, normalization, and simplification, and a (relative) lack of lexical diversity Teich (2003); Volansky et al. (2013); Toral (2019).
We test the approach on 20 translation directions, experimenting with 3 multilingual NMT models to predict the translation probabilities of human translations, NMT-produced translations, and pre-neural translations. We find that the approach detects the translation direction of human translations with an accuracy of 66% on average on the sentence level, and 80% for documents with 10 sentences. For the output of neural MT systems, detection accuracy is even higher, but our hypothesis that does not hold for the output of pre-neural systems.
Finally, we apply our method to a recent forensic case (Figure 2), where the translation direction of a German PhD thesis and an English book has been under dispute, finding additional evidence for the hypothesis that the English book is a forgery created to make the thesis appear plagiarized.
Our main contributions are the following:
We propose a simple, unsupervised approach to translation direction detection based on the translation probabilities of NMT models.
We demonstrate that the approach is effective for detecting the original translation direction of neural translations, and to a lesser extent, human translations in a variety of high-resource language pairs.
We provide a qualitative analysis of detection performance and apply the method to a real-world forensic case.
@@ 3 Methods
Given a parallel sentence pair , the main task in this work is to identify the translation direction between a language X and a language Y, and, consequently, establish which side is the original and which is the translation. This is achieved by comparing the conditional translation probability by an NMT model with the conditional translation probability by a model operating in the inverse direction. Our core assumption is that translations are assigned higher conditional probabilities than the original by NMT models, so if , we predict that is the translation, and the original, and the original translation direction is .
@@ 3.1 Detection on the Sentence Level
We follow earlier work by Junczys-Dowmunt (2018); Thompson and Post (2020), and average token-level (log-)probabilities.333 The models we use have been trained with label smoothing Szegedy et al. (2016), which has a cumulative effect on sequence-level probabilities Yan et al. (2023). Averaging token-level probabilities can help mitigate this shortcoming.
@@ 4 Experiments: Models and Data
We experiment with three massively multilingual machine translation models: M2M-100-418M (Fan et al., 2021), SMaLL-100 (Mohammadshahi et al., 2022), and NLLB-200-1.3B (NLLB Team et al., 2022).
@@
Total |
As an additional “real-world” dataset, we use 86 parallel sentences in German and English from a publicly documented plagiarism allegation case, in which translation-based plagiarism was the main focus (Ebbinghaus, 2022; Zenthöfer, 2022). The sentences were extracted from aligned excerpts of both the PhD thesis and the alleged source that were presented in a plagiarism analysis report (Weber, 2022). We extracted the sentences with OCR and manually checked for OCR errors. This dataset introduces an element of real-world complexity, as it involves translations that might not adhere strictly to professional standards nor does it have to fit into either of the aforementioned categories of translation strategy. It also provides a unique opportunity to test the robustness and adaptability of the translation direction detection system in a scenario that extends beyond controlled environments.
@@ 5.1 Sentence-level Classification
A second result is that the translation detection works best for neural translations (75.0% macro-average), second-best for human translations (66.5% macro-average), and worst for pre-neural (41.5% macro-average). The fact that performance for pre-neural systems is below chance level indicates that the NMT systems we use tend to assign low probabilities to the (often ungrammatical) outputs of pre-neural systems. In practice, one could pair our detection method with a monolingual model to identify such low-quality outputs.
@@
Language Pair | Ratio of | Ratio of |
HT bnhi | 66.80 % | 33.20 % |
HT csuk | 42.69 % | 57.31 % |
HT defr | 84.78 % | 15.22 % |
HT xhzu | 45.06 % | 54.94 % |
Sentences | Rel. Difference | |||
1 | DE: Mit dem Programm "Guten Tag, liebes Glück" ist er seit 2020 auf Tour. | |||
EN: He has been on tour with the programme "Guten Tag, liebes Glück" since 2020. (HT) | 0.145 | 0.558 | 0.26 | |
EN: He has been on tour since 2020. (NMT) | 0.272 | 0.092 | 2.95 | |
2 | EN: please try to perfprm thsi procedures" | |||
DE: bitte versuchen Sie es mit diesen Verfahren (HT) | 0.246 | 0.010 | 24.29 | |
DE: Bitte versuchen Sie, diese Prozeduren durchzuführen" (NMT) | 0.586 | 0.025 | 23.59 | |
3 | EN: If costs for your country are not listed, please contact us for a quote. | |||
DE: Wenn die Kosten für Ihr Land nicht aufgeführt sind, wenden Sie sich für einen Kostenvoranschlag an uns. (HT) | 0.405 | 0.525 | 0.77 | |
DE: Wenn die Kosten für Ihr Land nicht aufgeführt sind, kontaktieren Sie uns bitte für ein Angebot. (NMT) | 0.697 | 0.585 | 1.19 | |
4 | EN: Needless to say, it was chaos. | |||
DE: Es war natürlich ein Chaos. (HT) | 0.119 | 0.372 | 0.32 | |
DE: Unnötig zu sagen, es war Chaos. (NMT) | 0.755 | 0.591 | 1.28 | |
5 | DE: Mit freundlichen Grüßen | |||
FR: Cordialement (HT) | 0.026 | 0.107 | 0.24 | |
FR: Sincèrement (NMT) | 0.015 | 0.083 | 0.18 | |
FR: Sincères amitiés (NMT) | 0.062 | 0.160 | 0.39 | |
FR: Avec mes meilleures salutations (NMT) | 0.215 | 0.353 | 0.61 |
@@ 5.4 Qualitative Analysis
Misclassified short sentences as in Example 5 are not a rarity in our experiments. Our findings show that reliable detection of translation direction, with an average accuracy exceeding 50%, is consistently attained for all language pairs we tested starting at a sentence length between 50 and 60 characters.555We used SMaLL-100 for this analysis. Additionally, we observed a trend where the accuracy of direction detection increases as the length of the sentences grows. This aligns with previous unsupervised approaches, which also documented higher accuracy the larger the text chunks that were used, although there, reliable results were reported on a more extreme scale starting from text chunks with a length of 250 tokens Rabinovich and Wintner (2015).
5.5 Document-Level Classification
Accuracy scores for document-level results by M2M-100 (best-performing system at sentence level) are presented in Tables 8 (NMT) and 7 (HT). We consider documents with at least 10 sentences, and language pairs with at least 100 such documents in both directions.
@@ 5.5 Document-Level Classification
5.6 Application to Real-World Forensic Case
@@
Overall, our analysis supports the hypothesis that German is indeed the language of origin in this real-world dataset (forgery hypothesis; Figure 2). Nevertheless, we recommend that additional evidence be considered before drawing a final conclusion, given the error rate of 5–21% that we observed in our earlier experiments on English–German WMT documents.
@@ 6 Conclusion
We proposed a novel approach to detecting the translation direction of parallel sentences, using only an off-the-shelf multilingual NMT system. Experiments on WMT data showed that our approach, without any task-specific supervision, is able to detect the translation direction of NMT-translated sentences with relatively high accuracy. The accuracy increases to 96% if the classifier is provided with at least 10 sentences per document. We also found a robust accuracy for translations by human translators. Finally, we applied our approach to a real-world forensic case and found that it supports the hypothesis that the English book is a forgery.
@@
Low-resource languages:
Our experiments required test data for both translation directions, which limited the set of languages we could test. While the community has created reference translations for many low-resource languages, the translation directions are usually not covered symmetrically. For example, the test set of FLORES Goyal et al. (2022) has been translated from English into many languages, but not vice versa. Thus, apart from Table 5, we have not tested our approach on low-resource languages, and it is possible that the accuracy of our approach is lower for such languages, in parallel with the lower translation quality of NMT models for low-resource languages.
@@ Appendix B Comparison of Models (Document Level)
Appendix C Example for Forensic Dataset
DEEN | ENDE | |
EN: After 30 sec, the trypsin-containing PBS was poured off, and the cells were placed in the incubator for another minute. | 0.442 | 0.285 |
Appendix D Data Statistics